# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v2

@@ II.2 Simulation with circuit networks
$\mathcal{L}\rightarrow\mathcal{L}+\frac{1}{2L_{j}}\left(\frac{\phi_{j+1}-\phi_{j}}{2e}\right)^{2}\ .$ (8)
@@
However, this alone does not provide a tilt of the dispersion relation. This is where gyrators come into play. As pointed out in Ref. [71], a three-port gyrator, connecting two neighbouring nodes $\phi_{j+1}$ and $\phi_{j}$ to ground (whose phase is set to zero) can be described by the following contribution to the Lagrangian,
@@
where the parameter $G$ is proportional to the transconductance of the gyrator,
@@
At any rate, some of the lattice-specific stability issues can be appreciated on the general level here, others will be shown subsequently for a specific model. On a general footing, we see in Eq. (2) that any (constant or spatially dependent) local tilt $v$ (which in our simulation would be implemented by constant or spatially varying gyrator parameters $G_{j}$) must equally appear in the $(\partial_{x}\phi)^{2}$ term, as the corresponding matrix element $g^{11}$ ($\mu,\nu=1$ is the space-space component) is proportional to $v^{2}-u^{2}$. In other words, a finite nonreciprocity in the circuit needs to be “countered” by a corresponding inductive interaction, in order to realize actions of the form of Eq. (2). Consequently, we would need an according spatially dependent tuning of the local inductance profile $L_{j}$ which matches exactly with the values of $G_{j}$. But first of all, such a perfect fine tuning is realistically not possible, such that the condition $g=-1$ is in general only approximately, but never exactly, fulfilled. Hence, the mapping onto Eq. (5) is likewise not exact, and we cannot guarantee a perfect causal decoupling across the horizon, and the closed system might become spontaneously unstable over already short time scales. In fact, below we refrain from a simulation with constant metric determinant $g$ (in fact, we will modify the inductances rather than the gyrator values), such that the systems we consider do not map to Eq. (5) in the continuum limit. Furthermore, with the Hawking temperature being defined (at least in continuous field theories) as the derivative of the metric (i.e., the curvature) at the event horizon, there emerges the curious question of what happens to the observables of the system when considering a very sharp (discontinuous) change of the metric at the horizon, and how such a diverging curvature is regularized. In what follows we will provide an answer to that question for lattice simulators.
@@
    we find an important loop hole to the ambiguous black and white-hole nature of lattice event horizons (through an exceptional point in the dispersion relation) leading to a fundamentally different behaviour in the wormhole interior.
@@
$\begin{split}\sum_{j}G_{j}\left(\phi_{j}\dot{\phi}_{j+1}-\phi_{j+1}\dot{\phi}_{j}\right)=\quad\quad\\ \sum_{j}\dot{\phi}_{j}\left(G_{j-1}\phi_{j-1}-G_{j}\phi_{j+1}\right)\ ,\end{split}$ (13)
@@ III Setup
By means of standard circuit theory methods [62, 64], we find that the Hamiltonian of this array takes the form
$\mathcal{H}=$ $\sum_{j=1}^{J}E_{\text{C}}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}$
$-\sum_{j=1}^{J}E_{J}\cos(\phi_{j+1}-\phi_{j}+\phi_{\text{ext},j}),$ (14)
@@
FIG. 1: Wormhole simulations with quantum circuits. In circuits a) and b) (nearest and next-to-nearest Josephson coupling, respectively) the presence of gyrators tilts the dispersion relation giving left and right moving wave packets different speeds. The circuit c) is obtained by inverting the signs of a finite connected region inductances of circuit a). This is achieved by a flux quench induced by a localized current (inset), shifting the superconducting phase difference across the junction by $\pi$, leading to negative inductance. This creates two boundaries between the normal and over-tilted regions, which act as the apparent horizons. Horizons for the circuit with Josephson junctions connecting next-nearest neighbors b) can be obtained in a similar manner.
@@
$\mathcal{H}^{\prime}=$ $\sum_{j=1}^{J}E_{\text{C}}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}$
$-\sum_{j=1}^{J}E_{J}^{\prime}\cos(\phi_{j+1}-\phi_{j-1}+\phi_{\text{ext},j})\ .$ (15)
Considering the system at low energies (close to the many-body ground state), the phases on the nodes will approach an equilibrium configuration minimizing the total Josephson junction energy of the array. For sufficiently long arrays, we find either $\phi_{j+1}-\phi_{j}\approx-\phi_{\text{ext},j}+\delta\phi_{j+1}-\delta\phi_{j}$ (for $\mathcal{H}$) or $\phi_{j+1}-\phi_{j-1}\approx-\phi_{\text{ext},j}+\delta\phi_{j+1}-\delta\phi_{j-1}$ (for $\mathcal{H}^{\prime}$). Assuming $E_{C}<E_{J}$, the quantum fluctuations around the equilibrium value will be small, $\delta\phi_{j}\ll 1$. Consequently, the Josephson term in Hamiltonians (III) and (III) can be approximated by
$\approx\sum_{j=1}^{J}E_{L}(\delta\phi_{j+1}-\delta\phi_{j})^{2}+\text{const.}$.
$\approx\sum_{j=1}^{J}E_{L}^{\prime}(\delta\phi_{j+1}-\delta\phi_{j-1})^{2}+\text{const.},$. , (16)
@@
With the quadradic approximation of the Josephson energies, the total Hamiltonian describes a noninteracting boson field. Moreover, close to the ground state, the externally applied flux $\phi_{\text{ext},j}$ could be eliminated from the problem. It would therefore seem that neither the nonlinearity of the Josephson junction, nor the external flux play a role. However, as we will show in Sec. IV, transient flux-control and the compact cosine behaviour of the Josephson energy will allow us to generate nontrivial features related to quantum gravity. In particular, we will be able to modulate the effective inductive energy $E_{L}$ highly locally, and in particular, render it negative, allowing for the creation of instabilities that will lead to non-thermal Hawking radiation. Note that the nonlinearity of the Josephson energy was likewise of importance for earlier proposals to measure Hawking-like radiation emerging from soliton excitations in Josephson junction arrays [33]. We will comment on similarities with – but also significant differences to – our approach further below.
For the above conventional non-interacting boson problem, translational invariance allows us to analytically diagonalize the Hamiltonians (III) and (III), $\mathcal{H}^{(\prime)}=\sum_{m}\omega_{m}^{(\prime)}b_{m}^{\dagger}b_{m}$ (where index $m$ goes from $-J/2+1$ to $J/2$), to get the dispersion relations (Appendix A)
@@
where $b_{m}$ ($b_{m}^{\dagger}$) are bosonic annihilation (creation) operators satifying $[b_{m},b_{m^{\prime}}^{\dagger}]=\delta_{mm^{\prime}}$ and
@@
$\beta_{m}^{\prime}$ $=\frac{4E_{C}G^{2}\sin^{2}\left(\frac{2\pi m}{J}\right)+4E^{\prime}_{L}\sin^{2}\left(\frac{2\pi m}{J}\right)+M}{E_{C}},$ (18)
@@
For small $k$, the system exhibits a linear dispersion relation, [see Figs. 2a) and 2b)],
@@
In the relativistic sense, this would correspond to the analog speed of light, where the two branches $k>0$ and $k<0$ denote right and left moving signals, respectively [see also Eq. (6)]. Importantly, we see in general different analog speeds of light  $u+v$ ($u-v$) for excitations moving to the left, $k<0$ (right, $k>0$) [as illustrated in Figure 1a) and b)], with $u\propto 2\sqrt{4E_{C}^{2}G^{2}+E_{C}E_{L}}$ and $v\propto 4E_{\text{C}}G$. A nonzero $v$ corresponds to a tilt of the dispersion relation, which is only possible for a nonzero gyrator, $G\neq 0$, due to the aforementioned “conveyer belt” dynamics.
@@
FIG. 2: The above two figures show that the low-momentum eigenvalues (blue lines) of the circuit Hamiltonians (III) and (III) can be approximated by the linear dispersion relation (19) (orange lines) at low energies. The branches of the dispersion relation correspond to right and left movers with speeds $u\pm v$. One crucial point about the eigenvalues in b) is that near $k=\pi$ the dispersion relation (brown lines) is the mirror image of the dispersion near $k=0$, which is a feature that the eigenvalues in a) do not have. Parameters used: $E_{\text{C}}/E_{L}=1.3,E^{\prime}_{L}/E_{L}=0.25,G=0.6$, and $J=50$. To lift the degeneracy at zero energy, a small mass term is used $(\approx 10^{-3}E_{L}$).
Moreover, we observe that on this coarse-grained level (small $k$ essentially corresponds to taking the continuum limit of the lattice grid), the two models exhibit the exact same dispersion relation. Note however, that the next-nearest neighbour model hosts a second low-energy point with linear dispersion relation at $k\approx\pi$ [see Fig. 2b)]. The dispersion tilt here is always opposite to the one at $k\approx 0$. This feature emerges for the following reason. In the absence of the gyrators, the model $\mathcal{H}^{\prime}$ has two disjoint low-energy fields for even and odd lattice sites, resulting in a $\pi$-periodic Brillouin zone. The gyrators couple these two separate modes, restoring $2\pi$-periodicity with respect to $k$ (due to the alternating tilt), but keeping the two low-energy solutions intact. In essence, while model $\mathcal{H}$ hosts a single low-energy analog light field, model $\mathcal{H}^{\prime}$ hosts two separate fields with opposite dispersion relation. This detail will play an even more prominent role in what follows.
@@
FIG. 3: The principle behind the topological loop-hole for black versus white hole horizons in lattice systems via exceptional point. In panels a) and b), we show the complete dispersion of the circuit Hamiltonians (III) and (III) with negative inductances. When the Josephson junctions connect nearest neighbors, a), the eigenspectrum is complex (blue line represents real part of the eigenvalues and yellow line the imaginary part). The points where the eigenvalues go from real to complex are exceptional points (marked by red dots). Consequently, the spectrum only crosses zero once at $k=0$, and satisfies the periodicity constraint of the Brillouin zone via a detour in the complex plane. For next nearest neighbor coupling, b), the spectrum is real for all values of $k$, and crosses zero twice, again with mirror images at $k=0$ and $k=\pi$. The low momentum eigenvalues (orange line) for both the circuits do show an overtilt, which is required to create a horizon. Parameters are same as in Fig. 2, except the signs for $E_{L}$ and $E_{L}^{\prime}$ are flipped.
In the remainder of this section, we proceed by discussing the consequences a negative inductance has on the dispersion relation. As already stated, for small $k$ we thus get an overtilted linear dispersion relation in both models, see Figs. 3a) and 3b). As the slope of the dispersion relation now has the same sign for both positive and negative $k$, we get signals propagating in the same direction. If we now join regions with regular ($u>|v|$) and overtilted ($u<|v|$) dispersion relation, we could indeed engineer analog wormholes. But before we embark on that, we need to discuss in detail the behaviour of the dispersion relations, Eq. (III), for large $k$. Namely, when considering the full Brillouin zone (i.e., for $-\pi<k<\pi$), we see that the two models (nearest versus next-nearest neighbour) differ significantly. In particular, for the nearest neighbor coupling, we observe that outside the range $-k_{0}<k<k_{0}$ the dispersion relation becomes complex, see Fig. 3a), where $k_{0}$ is determined by the condition
@@
This transition point is equivalent to an exceptional point in the system. Crucially, this finding is in stark contrast to the model circuit where the Josephson junctions connect the next nearest neighbors. Here, for parameter regime $04E_{C}G^{2}+E_{L}>$ 0, we can overtilt the dispersion relation without encountering complex eigenvalues, see Fig. 3b).
@@
There are several reasons why both of these models are interesting in their own right. First, let us comment in more detail on stability. In principle, the inversion of inductance (or capacitance for that matter) indicates a spontaneous electrostatic instability. This can be understood already on a much simpler level of a single $LC$ circuit, with a resonance frequency $\sim\sqrt{E_{L}E_{C}}$. Swapping the sign of the inductance leads to an imaginary frequency. Note that for transient times, this result is by no means problematic (see also Sec. II.1), as such an imaginary frequency corresponds to placing a localized wave function on the tip of an inverted harmonic potential. For transient times, we can by all means obtain physically meaningful results from the Schrödinger equation. In particular, note that such an evolution will result in an exponential blowing up of quantum fluctuations of both canonically conjugate quantities, here, charge and phase (as we will see later also in explicit calculations). This might at first sight seem counterintuitive. The most common situation for generic quantum wave functions is that the uncertainty in one of the two conjugate spaces is inversely proportional to the uncertainty in the other. But note that inverted potentials provide very special wave functions over time that spread both in position (e.g., $\phi$) space, as the wave function evolves from the tip of the parabola to both sides, as well as increasing uncertainty in momentum (e.g., $N$) space, since the particle wave gets more and more accelerated as it rolls down both sides of the inverted parabola. The theory is only problematic in the long-time limit, due to the formal removal of a well-defined ground state. We note that our realization of negative inductances via the inherent nonlinearity of the Josephson effect provides a very neat conceptual realization of instabilities, while (in principle) retaining a well-defined ground state for long times, as we explain in more detail further below. We also note that a recent work [104] has explored a related but distinct concept of negative mass resonators in cQED architectures, by strongly driving a weakly nonlinear superconducting LC circuit [105], leading effectively to negative susceptibility.
@@
To proceed, we note that in the 1D chain models we consider, there are ways to stabilize the system in spite of negative inductances, in line with the discussion in Sec. II. Indeed, the gyrators play the pivotal role for stabilization: in the Hamiltonians of Eqs. (III) and (III), when we expand the term in the first line $\sim E_{C}$, we get an effective inductive element due to a nonzero $G$, $\sim E_{C}G^{2}(\phi_{j+1}-\phi_{j-1})^{2}$. This is an effective inductive contribution due to the gyrator, which is guaranteed to be positive even when $E_{L}<0$ (as long as $E_{C}>0$). But as already foreshadowed in Sec. II.2, this effective inductive contribution couples next-nearest neighbour sites. Hence, a negative next-nearest inductor [Eq. (III)] can be exactly compensated by a positive next-nearest neighbour contribution of the gyrator. The negative nearest neighbour inductor model [Eq. (III)] on the other hand cannot exactly be compensated by the gyrators, hence the presence of the exceptional point.
@@
Let us now take a bit of a more formal perspective on the above observation for generic wormhole simulations, which will allow us to demonstrate one of the central results of this work, see also item (iv) in Sec. II.2. Namely, there is a simple topological connection between overtilted spectra, Brillouin zone sizes, and exceptional points. For discrete lattice implementations the dispersion relation is always $2\pi$ periodic in $k$. Crucially, an overtilted spectrum means that the energies cross from positive to negative values at the transition points $k=0,2\pi,\ldots$. If the spectrum is real for all $k$, then periodicity in momentum space always guarantees that an overtilted spectrum near a certain value of $k$ must have an overtilted partner at another value of $k$ with opposite tilt direction, as a periodic spectrum must traverse the zero-energy line an even number of times. Consequently, for any lattice implementation, it seems unavoidable that what appears like a black hole horizon near one value of $k$, must necessarily have a white hole horizon partner at another value of $k$. This fact has already been pointed out in the context of analog black-hole simulations in tilted Weyl semimetal structures [31]. Our next-nearest neighbour model confirms this expectation, where the overtilt point at $k\approx 0$ has an inverted partner point at $k\approx\pm\pi$, see Fig. 3 b). These two $k$ points can be individually addressed when preparing, e.g., Gaussian wave packets centered either around $k\approx 0$ (wave packets that are smooth in $j$ space) and $k\approx\pm\pi$ (wave packets with alternating even-odd signs in $j$ space).

### v4

@@ 2.2 Simulation with circuit networks
$\mathcal{L}\rightarrow\mathcal{L}-\frac{1}{2L_{j}}\left(\frac{\phi_{j+1}-\phi_{j}}{2e}\right)^{2}\ .$ (8)
@@
However, this alone does not provide a tilt of the dispersion relation. This is where gyrators come into play, see Fig. 1 for a schematic representation and a possible realization of a mesoscopic quantum gyrator, recently developed in Ref. [72]. As pointed out in Ref. [67], a three-port gyrator, connecting two neighbouring nodes $\phi_{j+1}$ and $\phi_{j}$ to ground (whose phase is set to zero) can be described by the following contribution to the Lagrangian,
@@
where $G$ parametrizes the transconductance of the gyrator. Note that the gyrator Lagrangian can be subject to gauge transformations, such that its form is not uniquely determined. In the above chosen symmetric gauge [72], it can be directly related to the transport properties across the device,
@@
[…]
Figure 1: a) Schematic of a gyrator, a four-terminal circuit element, introduced by Tellegen [65]. Here two of its terminals have been grounded, hence it acts as an effective two terminal element. b) One possible realization of a (quantum) gyrator, as proposed in [72], by using quantum dots proximitized with superconductors. Building up on the prediction of Ref. [73] that conventional multiterminal Josephson junctions harbour topological bands in the space of superconducting phases, one of the goals of Ref. [72] was the development of blueprints to create topological flat bands, where the circuit behaviour reduces to Eqs. (10) and (11). The above quantum dot array is one of three proposed ways to generate topological flat bands in multiterminal junctions. Flat bands emerge under appropriate tuning of the central flux ($\phi_{0}$) and for sufficiently long chains ($J$ large). For further details, see Ref. [72].
@@
At any rate, some of the lattice-specific stability issues can be appreciated on the general level here, others will be shown subsequently for a specific model. On a general footing, we see in Eq. (2) that any (constant or spatially dependent) local tilt $v$ (which in our simulation would be implemented by constant or spatially varying gyrator parameters $G_{j}$) must equally appear in the $(\partial_{x}\phi)^{2}$ term, as the corresponding matrix element $g^{11}$ ($\mu,\nu=1$ is the space-space component) is proportional to $v^{2}-u^{2}$. In other words, a finite nonreciprocity in the circuit needs to be “countered” by a corresponding inductive interaction, in order to realize actions of the form of Eq. (2). Consequently, we would need an according spatially dependent tuning of the local inductance profile $L_{j}$ which matches exactly with the values of $G_{j}$. But first of all, such a perfect fine tuning is realistically not possible, such that the condition g=1g=-1 is in general only approximately, but never exactly, fulfilled. Hence, the mapping onto Eq. (5) is likewise not exact, and we cannot guarantee a perfect causal decoupling across the horizon, and the closed system might become spontaneously unstable over already short time scales. In fact, below we refrain from a simulation with constant metric determinant $g$ (in fact, we will modify the inductances rather than the gyrator values), such that the systems we consider do not map to Eq. (5) in the continuum limit. Furthermore, with the Hawking temperature being defined (at least in continuous field theories) as the derivative of the group velocity (i.e., the surface gravity) at the event horizon, there emerges the curious question of what happens to the observables of the system when considering a very sharp (discontinuous) change of the metric at the horizon, and how such a diverging surface gravity is regularized. In what follows we will provide an answer to that question for lattice simulators. It should be noted that step like horizons have been studied in continuum approximation in the context of BECs [123, 124]. In these works, unlike ours, scattering modes with real eigenenergies form a complete basis. Therefore, it is still possible to define a scattering mode where an incoming negative norm mode scatters to an outgoing positive norm mode. This mode is often termed the "Hawking mode", and its flux can be used to deduce the spectrum of the Hawking radiation.
@@
    we find an important loop hole to the ambiguous black and white-hole nature of lattice event horizons (through an EP in the dispersion relation) leading to a fundamentally different behaviour in the wormhole interior.
@@
$\sum_{j}G_{j}\left(\phi_{j}\dot{\phi}_{j+1}-\phi_{j+1}\dot{\phi}_{j}\right)=\sum_{j}\dot{\phi}_{j}\left(G_{j-1}\phi_{j-1}-G_{j}\phi_{j+1}\right)\ ,$ (13)
@@ 3 Setup
The Lagrangian of this circuit can be written as
$\mathcal{L}$ $=\sum_{j=1}^{J}\left[\frac{C}{2}\left(\frac{\dot{\phi}_{j}}{2e}\right)^{2}+G\phi_{j}\left(\dot{\phi}_{j+1}-\dot{\phi}_{j-1}\right)+E_{J}\cos\left(\phi_{j+1}-\phi_{j}+\phi_{\text{ext},j}\right)\right].$ (14)
To get to the Hamiltonian we first define the conjugate variable $N_{j}=\partial\mathcal{L}/\partial\dot{\phi}_{j}$ and then perform the Legendre transform $\mathcal{H}=\sum_{j}\dot{\phi}_{j}N_{j}$-$\mathcal{L}$, in alignment with standard circuit quantization [53, 55]. We find that the Hamiltonian of this array takes the form
$\mathcal{H}=$ $\sum_{j=1}^{J}\left[E_{\text{C}}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}-E_{J}\cos(\phi_{j+1}-\phi_{j}+\phi_{\text{ext},j})\right],$ (15)
@@
Figure 2: Wormhole simulations with quantum circuits. In circuits a) and b) (nearest and next-to-nearest Josephson coupling, respectively) the presence of gyrators tilts the dispersion relation giving left and right moving wave packets different speeds. The circuit c) is obtained by inverting the signs of a finite connected region inductances of circuit a). This is achieved by a flux quench induced by a localized current (inset), shifting the superconducting phase difference across the junction by $\pi$, leading to negative inductance. This creates two boundaries between the normal and over-tilted regions, which act as the apparent horizons. Horizons for the circuit with Josephson junctions connecting next-nearest neighbors b) can be obtained in a similar manner.
@@
$\mathcal{H}^{\prime}=$ $\sum_{j=1}^{J}\left[E_{\text{C}}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}-E_{J}^{\prime}\cos(\phi_{j+1}-\phi_{j-1}+\phi_{\text{ext},j})\right]\ .$ (16)
Considering the system at low energies (close to the many-body ground state), the phases on the nodes will approach an equilibrium configuration minimizing the total Josephson junction energy of the array. For sufficiently long arrays, we find either $\phi_{j+1}-\phi_{j}\approx-\phi_{\text{ext},j}+\delta\phi_{j+1}-\delta\phi_{j}$ (for $\mathcal{H}$) or $\phi_{j+1}-\phi_{j-1}\approx-\phi_{\text{ext},j}+\delta\phi_{j+1}-\delta\phi_{j-1}$ (for $\mathcal{H}^{\prime}$). Assuming $E_{C}<E_{J}$, the quantum fluctuations around the equilibrium value will be small, $\delta\phi_{j}\ll 1$. Consequently, the Josephson term in Hamiltonians (15) and (16) can be approximated by
$\approx\sum_{j=1}^{J}E_{L}(\delta\phi_{j+1}-\delta\phi_{j})^{2}+\text{const.}$
$\approx\sum_{j=1}^{J}E_{L}^{\prime}(\delta\phi_{j+1}-\delta\phi_{j-1})^{2}+\text{const.},$ (17)
@@
With the quadradic approximation of the Josephson energies, the total Hamiltonian describes a noninteracting boson field. Moreover, close to the ground state, the externally applied flux $\phi_{\text{ext},j}$ could be eliminated from the problem. It would therefore seem that neither the nonlinearity of the Josephson junction, nor the external flux play a role. However, as we will show in Sec. 4, transient flux-control and the compact cosine behaviour of the Josephson energy will allow us to generate nontrivial features related to quantum gravity. In particular, we will be able to modulate the effective inductive energy $E_{L}$ highly locally, and in particular, render it negative, allowing for the creation of instabilities that will lead to non-thermal Hawking radiation. Note that the nonlinearity of the Josephson energy was likewise of importance for earlier proposals to measure Hawking-like radiation emerging from soliton excitations in Josephson junction arrays [33]. We will comment on similarities with – but also significant differences to – our approach further below.
For the above conventional non-interacting boson problem, translational invariance allows us to analytically diagonalize the Hamiltonians (15) and (16), $\mathcal{H}^{(\prime)}=\sum_{m}\omega_{m}^{(\prime)}b_{m}^{\dagger}b_{m}$ (where index $m$ goes from J/2+1-J/2+1 to $J/2$), to get the dispersion relations (Appendix A)
@@
where $b_{m}$ ($b_{m}^{\dagger}$) are bosonic annihilation (creation) operators satifying $b_{m},b_{m^{\prime}}^{\dagger}]=\delta_{mm^{\prime}}$ and
@@
$\beta_{m}^{\prime}$ $=\frac{4E_{C}G^{2}\sin^{2}\left(\frac{2\pi m}{J}\right)+4E^{\prime}_{L}\sin^{2}\left(\frac{2\pi m}{J}\right)+M}{E_{C}},\$(19)
@@
For small $k$, the system exhibits a linear dispersion relation, [see Figs. 3a) and 3b)],
@@
In the relativistic sense, this would correspond to the analog speed of light, where the two branches $k>0$ and $k<0$ denote right and left moving signals, respectively [see also Eq. (6)]. Importantly, we see in general different analog speeds of light  $u+v$ (uvu-v) for excitations moving to the left, $k<0$ (right, $k>0$) [as illustrated in Figure 2a) and b)], with $u\propto 2\sqrt{4E_{C}^{2}G^{2}+E_{C}E_{L}}$ and $v\propto 4E_{\text{C}}G$. A nonzero $v$ corresponds to a tilt of the dispersion relation, which is only possible for a nonzero gyrator, $G\neq 0$, due to the aforementioned “conveyer belt” dynamics.
@@
Figure 3: The above two figures show that the low-momentum eigenvalues (blue lines) of the circuit Hamiltonians (15) and (16) can be approximated by the linear dispersion relation (20) (orange lines) at low energies. The branches of the dispersion relation correspond to right and left movers with speeds $u\pm v$. One crucial point about the eigenvalues in b) is that near $k=\pi$ the dispersion relation (brown lines) is the mirror image of the dispersion near $k=0$, which is a feature that the eigenvalues in a) do not have. Parameters used: $E_{\text{C}}/E_{L}=1.3,E^{\prime}_{L}/E_{L}=0.25,G=0.6$, and $J=50$. To lift the degeneracy at zero energy, a small mass term is used $(\approx 10^{-3}E_{L}$).
Moreover, we observe that on this coarse-grained level (small $k$ essentially corresponds to taking the continuum limit of the lattice grid), the two models exhibit the exact same dispersion relation. Note however, that the next-nearest neighbour model hosts a second low-energy point with linear dispersion relation at $k\approx\pi$ [see Fig. 3b)], in accordance with Ref. [48]. The dispersion tilt here is always opposite to the one at $k\approx 0$. This feature emerges for the following reason. In the absence of the gyrators, the model $\mathcal{H}^{\prime}$ has two disjoint low-energy fields for even and odd lattice sites, resulting in a $\pi$-periodic Brillouin zone. The gyrators couple these two separate modes, restoring $2\pi$-periodicity with respect to $k$ (due to the alternating tilt), but keeping the two low-energy solutions intact. In essence, while model $\mathcal{H}$ hosts a single low-energy analog light field, model $\mathcal{H}^{\prime}$ hosts two separate fields with opposite dispersion relation. This detail will play an even more prominent role in what follows.
@@
Figure 4: The principle behind the topological loop-hole for black versus white hole horizons in lattice systems via EPs. In panels a) and b), we show the complete dispersion of the circuit Hamiltonians (15) and (16) with negative inductances. When the Josephson junctions connect nearest neighbors, a), the eigenspectrum is complex (blue line represents real part of the eigenvalues and yellow line the imaginary part). The points where the eigenvalues go from real to complex are EPs (marked by red dots). Consequently, the spectrum only crosses zero once at $k=0$, and satisfies the periodicity constraint of the Brillouin zone via a detour in the complex plane. For next nearest neighbor coupling, b), the spectrum is real for all values of $k$, and crosses zero twice, again with mirror images at $k=0$ and $k=\pi$. The low momentum eigenvalues (orange line) for both the circuits do show an overtilt, which is required to create a horizon. Parameters are same as in Fig. 3, except the signs for $E_{L}$ and $E_{L}^{\prime}$ are flipped.
In the remainder of this section, we proceed by discussing the consequences a negative inductance has on the dispersion relation. As already stated, for small $k$ we thus get an overtilted linear dispersion relation in both models, see Figs. 4a) and 4b). As the slope of the dispersion relation now has the same sign for both positive and negative $k$, we get signals propagating in the same direction. If we now join regions with regular ($u>|v|$) and overtilted ($u<|v|$) dispersion relation, we could indeed engineer analog wormholes. But before we embark on that, we need to discuss in detail the behaviour of the dispersion relations, Eq. (3), for large $k$. Namely, when considering the full Brillouin zone (i.e., for $\pi<k<\pi$), we see that the two models (nearest versus next-nearest neighbour) differ significantly. In particular, for the nearest neighbor coupling, we observe that outside the range $k_{0}<k<k_{0}$ the dispersion relation becomes complex, see Fig. 4a), where $k_{0}$ is determined by the condition
@@
This transition point is equivalent to an EP in the system. Crucially, this finding is in stark contrast to the model circuit where the Josephson junctions connect the next nearest neighbors. Here, for parameter regime $4E_{C}G^{2}+E_{L}>0$, we can overtilt the dispersion relation without encountering complex eigenvalues, see Fig. 4b).
Before moving ahead we would like to address the seeming incongruity of having complex eigenvalues for a Hermitian Hamiltonian, as we just observed in the case of nearest neighbor coupling with negative inductance. We did preempt this in Sec.2.1, by stating that existence of complex eigenvalues implies instability in the system. But still this result might be confusing as the Hamiltonian we are working with remains Hermitian at all times, in short this problem comes around because we are being cavalier with the word "eigenvalues". Looking at Appendices A and B, one can see that to diagonalize the Hamiltonian we have to eventually perform a bosonic Bogoliubov transformation. It is a well known fact that for Bosons the Bogoliubov transformation involves finding eigenvalues of a non-Hermitian matrix. If these eigenvalues are real then one can go ahead and diagonalize the original second quantized Hamiltonian, and realize that the eigenvalues of the second quantized Hamiltonian correspond to the eigenvalues of the non-Hermitian matrix. On the other hand, if the eigenvalues of the non-Hermitian matrix are complex, the idea that Bogoliubov transformation is diagonalizing the second quantized Hamiltonian goes out the window [122]. Instead the Bogoliubov transformation brings the Hamiltonian in the following form: $\mathcal{H}=\sum_{i}\lambda_{i}b^{\prime}_{i}b_{i}$. Here the operators $b^{\prime}_{i}$ and $b_{i}$ follow the CCR but are not related to each other via Hermitian conjugation, and $\lambda_{i}$, which are the eigenvalues of the non-Hermitian matrix, are complex. Still, this kind of Bogoliubov transformation is useful, as it makes studying the dynamics of the system much easier, which is what we do in Sec.5. A more thorough exploration of under what conditions a quadratic bosonic Hamiltonian is diagonalizable using Bogoliubov transformation can be found in [125].
@@
There are several reasons why both of these models are interesting in their own right. First, let us comment in more detail on stability. In principle, the inversion of inductance (or capacitance for that matter) indicates a spontaneous electrostatic instability. This can be understood already on a much simpler level of a single $LC$ circuit, with a resonance frequency $\sim\sqrt{E_{L}E_{C}}$. Swapping the sign of the inductance leads to an imaginary frequency. Note that for transient times, this result is by no means problematic (see also Sec. 2.1), as such an imaginary frequency corresponds to placing a localized wave function on the tip of an inverted harmonic potential. For transient times, we can by all means obtain physically meaningful results from the Schrödinger equation. In particular, note that such an evolution will result in an exponential blowing up of quantum fluctuations of both canonically conjugate quantities, here, charge and phase (as we will see later also in explicit calculations). This might at first sight seem counterintuitive. The most common situation for generic quantum wave functions is that the uncertainty in one of the two conjugate spaces is inversely proportional to the uncertainty in the other. But note that inverted potentials provide very special wave functions over time that spread both in position (e.g., $\phi$) space, as the wave function evolves from the tip of the parabola to both sides, as well as increasing uncertainty in momentum (e.g., $N$) space, since the particle wave gets more and more accelerated as it rolls down both sides of the inverted parabola. The theory is only problematic in the long-time limit, due to the formal removal of a well-defined ground state. We note that our realization of negative inductances via the inherent nonlinearity of the Josephson effect provides a very neat conceptual realization of instabilities, while (in principle) retaining a well-defined ground state for long times, as we explain in more detail further below. We also note that a recent work [126] has explored a related but distinct concept of negative mass resonators in cQED architectures, by strongly driving a weakly nonlinear superconducting LC circuit [127], leading effectively to negative susceptibility.
@@
To proceed, we note that in the 1D chain models we consider, there are ways to stabilize the system in spite of negative inductances, in line with the discussion in Sec. 2. Indeed, the gyrators play the pivotal role for stabilization: in the Hamiltonians of Eqs. (15) and (16), when we expand the term in the first line $\sim E_{C}$, we get an effective inductive element due to a nonzero $G$, $\sim E_{C}G^{2}(\phi_{j+1}-\phi_{j-1})^{2}$. This is an effective inductive contribution due to the gyrator, which is guaranteed to be positive even when $E_{L}<0$ (as long as $E_{C}>0$). But as already foreshadowed in Sec. 2.2, this effective inductive contribution couples next-nearest neighbour sites. Hence, a negative next-nearest inductor [Eq. (15)] can be exactly compensated by a positive next-nearest neighbour contribution of the gyrator. The negative nearest neighbour inductor model [Eq. (16)] on the other hand cannot exactly be compensated by the gyrators, hence the presence of the EP.
@@
Let us now take a bit of a more formal perspective on the above observation for generic wormhole simulations, which will allow us to demonstrate one of the central results of this work, see also item (iv) in Sec. 2.2. Namely, there is a simple topological connection between overtilted spectra, Brillouin zone sizes, and EPs. For discrete lattice implementations the dispersion relation is always $2\pi$ periodic in $k$. Crucially, an overtilted spectrum means that the energies cross from positive to negative values at the transition points $k=0,2\pi,\ldots$. If the spectrum is real for all $k$, then periodicity in momentum space always guarantees that an overtilted spectrum near a certain value of $k$ must have an overtilted partner at another value of $k$ with opposite tilt direction, as a periodic spectrum must traverse the zero-energy line an even number of times (essentially a simplified version of the Nielsen-Ninomiya theorem [48]). Consequently, for any lattice implementation, it seems unavoidable that what appears like a black hole horizon near one value of $k$, must necessarily have a white hole horizon partner at another value of $k$. This fact has already been pointed out in the context of analog black-hole simulations in tilted Weyl semimetal structures [31]. Our next-nearest neighbour model confirms this expectation, where the overtilt point at $k\approx 0$ has an inverted partner point at $k\approx\pm\pi$, see Fig. 4 b). These two $k$ points can be individually addressed when preparing, e.g., Gaussian wave packets centered either around $k\approx 0$ (wave packets that are smooth in $j$ space) and $k\approx\pm\pi$ (wave packets with alternating even-odd signs in $j$ space).